# LIGHTS. A robust technique to identify galaxy edges

### v1

@@ 3 Target selection and data
We apply the methodology described above to the LIGHTS data of NGC 3486, a M33-like galaxy (i.e., similar stellar and total mass to M33), located at 13.6 Mpc (Kourkchi et al. 2020) and in a region of low Galactic extinction (Ag = 0.08, Ar = 0.06;
@@
NGC 3486 constitutes a case study, we aim to perform the analysis to the entire LIGHTS dataset.
@@
LBCs setup includes four CCDs, with each CCD covering 7.8 17.6 section of the sky and a total Field of View (FoV) of the camera of 23 25.
@@
The data collection took place during 2021 April 3 under dark sky conditions. By using dithered exposures of 180 seconds each, with four arcminutes offsets, we obtained a total on-source integration time of 1.5 hours (see Zaritsky et al. 2024 for more details on observations).
The data reduction process followed a similar strategy than the one outlined in Trujillo et al. (2021) and developed in Zaritsky et al. (2024).
Our pipeline’s main steps involve deriving flat field images, background removal, astrometric and photometric calibrations, and mosaic co-addition.
@@
With this observational strategy and optimized data reduction we achieved a surface brightness limiting depth of 31.4 mag/arcsec2 and 30.5 mag/arcsec2 in and Sloan bands respectively, corresponding to a 3 fluctuation of the background in areas equivalent to 10 10 (Trujillo & Fliri 2016; Román et al. 2020).
@@ 4 Methods
The edge detection process for NGC 3486 follows a structured approach. We begin by masking foreground and background sources to minimize contamination. Next, although LIGHTS data are background-subtracted, in some cases, the subtraction was not precise enough to reliably represent the local surrounding background of the galaxy, leading to slight over- or underestimation. Therefore, we correct of the local background. Then, we apply the Point Spread Function (PSF) correction to account for the scattering of the galaxy’s light due to the effect of the PSF. After this, we proceed to get the surface stellar mass density map of NGC 3486 and finally we derive the location of the edge of NGC 3486 using the methodology described in Sect. 2. Figure 2 provides a schematic overview that serves as a reference for the methodology presented in the following sections.
@@ 4.1 Mask and local background treatment
To achieve accurate background subtraction, we first mask foreground and background sources. The mask is generated by combining the and band masks. The first step in creating the masks involves running NoiseChisel and Segment (Akhlaghi & Ichikawa 2015; Akhlaghi 2019). To improve the detection of contaminating sources, we apply unsharp masking, which removes the light from the galaxy and facilitates the identification of individual contamination sources. Unsharp masking involves smoothing the original image with a Gaussian filter corresponding to the FWHM of point-like sources and subtracting this smoothed image from the original. Finally, we manually inspect the deep color image of NGC 3486 (see Fig. 9) to mask any remaining undetected objects. The final mask applied to NGC 3486 is shown in Fig. 10 in Appendix C.
To compute the local background correction, once the data are masked, we manually place circular apertures (each with a 20 radius) in specific regions approximately 6.5 from the galaxy center. These locations are selected to be relatively free of contamination, avoiding bright sources ( mag) and diffuse light brighter than mag/arcsec2. We adopt this conservative approach to ensure robust background estimation. We then compute the resistant (3-clipped) median of the unmasked background pixels within these regions. This process is performed separately for each filter.
@@ 4.2 Removing the effect of the PSF in the outermost regions of the galaxy
Light scattering from the PSF can significantly alter the observed properties of a galaxy’s outer disc by redistributing mostly light from the central regions, leading to an artificial brightening of the outskirts (Trujillo et al. 2001; Sandin 2014, 2015). The effect depends on the instrument and wavelength used (Duc et al. 2015; Croston et al. 2006; Mihos et al. 2015; Abraham & van Dokkum 2014). In this work, we mitigate PSF effects in the galaxy’s outermost regions by combining model-based convolution techniques (Trujillo & Fliri 2016; Rich et al. 2019; Peters et al. 2017b; Martínez-Lombilla & Knapen 2019; Infante-Sainz et al. 2020) with direct image deconvolution methods, such as PyOperators (Karabal et al. 2017). The following sections detail the steps taken in this process.
@@ 4.2.1 Breaking the PSF-Stellar halo degeneracy
The problem: The presence of a faint stellar halo around galaxies can be challenging to distinguish from the light-broadening effects caused by the PSF. This confusion, that we term PSF-Stellar halo degeneracy, makes it difficult to tell if a galaxy’s outer light distribution is due to an extended stellar halo or merely an artifact of the PSF. To illustrate the problem, we generated two BrokenExponential models using the makeimage task in IMFIT, with the same PA, AR, and outer edge radius . Model1, has an outer slope () that is 20 shallower than the one of Model2. In addition, Model2 is composed also of a faint Exponential stellar halo component. We applied the extended band PSF of the LBC (Sedighi et al. 2025) to both models. The averaged stellar mass surface density of Model1 and Model2 are illustrated in the left and central panels of Fig. 3 respectively along with their components and the profile of the PSF. The parameters used to create the models are reported in Appendix D.
@@
The outer profile (R ¿ 200 pix) in both models can be interpreted as either only a PSF effect on the declining disc light or a faint stellar halo combined with the PSF. This uncertainty highlights the PSF-Stellar halo degeneracy, making it difficult to distinguish between PSF broadening and a different outer profile plus a faint stellar halo. This test can be extended to higher inclinations, which amplify PSF effects, emphasizing the need for accurate PSF deconvolution when analyzing faint outer structures in galaxies.
@@
In summary, the PSF redistributes photons, artificially broadening the outer regions of galaxies and smoothing the slope of the outer disc (). In general this depends on the shape of the PSF. This redistribution does not alter the location of the edge in the stellar mass density distribution.
However, it is key to correct for the PSF to ensure accurate recovery of the intrinsic shape of the profile, particularly the outer slope ().
@@
Wiener-Hunt deconvolution to break the degeneracy: We use the extended PSF of the LIGHTS survey (Sedighi et al. 2025) to deconvolve the background-subtracted data by applying the Wiener-Hunt (Orieux et al. 2010) deconvolution algorithm. To mathematically frame the effect of the PSF, the observed data from the telescope can be expressed as a convolution of the “true” image with the PSF, , plus an additive noise term :
@@
In an ideal, noise-free scenario, we can use Fourier transforms to work with these data components in frequency space. Specifically, the Fourier transforms of and of would allow us to isolate , the Fourier transform of the “true” image. From this, we could reconstruct by transforming back into spatial space. However, because acts as a low-pass filter, it tends to amplify high-frequency noise. Thus, a noise suppressor is needed. The deconvolution process using Wiener-Hunt filter is performed in the Fourier domain:
@@
where is the complex conjugate of and is the regularization parameter that controls noise suppression. In this work we use Wiener-Hunt deconvolution implemented in the Python package skimage666 skimage.restoration.unsupervised.Wiener, where the regularization parameter is automatically estimated using a stochastic iterative process (Orieux et al. 2010).
@@
To visualize the effect of performing the Wiener-Hunt deconvolution on the data, we extract the surface brightness profiles for NGC 3486 in both and bands with astscript-radial-profile and a bin size of 10777Larger bin sizes smooth the profiles by averaging the flux over wider elliptical annuli, thereby reducing point-to-point fluctuations. In contrast, smaller bin sizes increase the sensitivity to small-scale variations, leading to more oscillatory behavior in the derivative profiles. We verified that the choice of step size does not affect the determination of the edge location.. We extract the light profiles of the galaxy using AR = and a PA of (clockwise from the X-axis). This choice of parameters is justified in the Appendix E and compared with previous estimates. To maintain consistency, we apply the same mask used for sky background estimation (see Fig. 10). The errors on these profiles are calculated by combining the uncertainties of each elliptical annulus with those of the local sky estimate as outlined in Golini et al. (2024).
@@
In Fig. 4, we show in green the resulting and surface brightness profiles prior deconvolution, which exhibit an inner bulge component extending to about 40, followed by a exponential decline. At around 200, a slope change occurs, with the profile steepening in both bands, marking the edge of the visible disc seen in Fig. 9.
@@
The Wiener-Hunt deconvolved surface brightness profiles in the and bands are also shown in Fig. 4. Compared to the original (background subtracted) profiles, the deconvolved profiles exhibit a steeper outer slope beyond R = 190. We visually define this transition point as . This effect is due to the redistribution of light from the galaxy’s outskirts to the inner regions, compensating for the PSF-induced broadening in the original data. For instance, post-deconvolution, the inner region (R ¡ 20) appears 0.1 magnitudes brighter, while outer regions (R ¿ ″) are up to 1 magnitude fainter. The edge location is preserved.
@@
The Wiener-Hunt filter is a widely used deconvolution technique, but like other deconvolution methods such as Richardson-Lucy (Richardson 1972; Lucy 1974) or those based on spatial regularizations like the assumption of sparsity in a wavelet domain (Starck & Murtagh 1994; Starck et al. 2006), it introduces image smoothness, correlates pixel values, and alters background statistics, especially in regions of low surface brightness (LSB) (Murli et al. 1999; Orieux et al. 2010; Ramos et al. 2020; Yoo & Ahn 2014).
@@
4.2.2 Creation of the model of the galaxy not affected by PSF
To remove the PSF effect from the NGC 3486 data while preserving the noise structure of the LIGHTS image, we proceed as follows. First, we model the averaged surface brightness profile of NGC 3486 in both bands after Wiener-Hunt deconvolution (blue profiles in Fig. 4) using a custom implementation of the IMFIT Exponential function in Python.
We model the Wiener-Hunt deconvolved profile in the range between (vertical dashed purple lines in Fig. 4), which is the point where the PSF begins to significantly affect the outermost regions of the light profiles in the and filters (i.e., 25 mag/arcsec2), to the radius where the uncertainty falls below 0.1 mag. This results in an outer exponential scale length () of 18.8 in both filters. Second, we generate a two-dimensional image of this Exponential model using makeimage. When constructing the two-dimensional outer disc component, we adopt the PA and inclination parameters obtained in Appendix E. We do not include a bulge or stellar halo component to avoid imposing a priori assumptions. Stellar halos often exhibit asymmetries, spatial offsets, or orientations misaligned with the disc, making them difficult to model reliably. Our final image of the galaxy, not affected by the PSF, consists of two parts. An inner region (R ¡ ) represented by the Wiener-Hunt deconvolved image, which enhances spiral structures that would otherwise be difficult to model and an outer region (R ¿ ) which transitions to an exponential declining profile (shown in pink in Fig. 4).
@@ 4.2.3 Subtraction of the model and residuals addition: PSF deconvolved data
After generating the two deconvolved models (one for and one for the band) of NGC 3486, we employ the astconvolve tool in Gnuastro to convolve these models with the LIGHTS’ PSF.
To isolate structural components beyond the PSF effect, we subtract these convolved images from the original galaxy data using the astarithmetic tool in Gnuastro, resulting in a ”residuals” image. These residuals highlight the galaxy’s asymmetric faint features in the outer region of the galaxy that are not accounted for in the symmetric model.
Following this, we add the residuals image back into the model created in Sect. 4.2.2, resulting in a refined image with PSF effects effectively removed. This stage of analysis ensures that any faint, residual light detected in the galaxy’s outer regions corresponds to real, low-luminosity stellar structures rather than artifacts introduced by the PSF.
The surface brightness profiles, after PSF correction, extracted using fixed ellipses, are shown in orange in Fig. 4. In both bands, the galaxy light distribution after PSF removal shows a steeper decline, up to one magnitude fainter, starting at R ″. We remark that, to avoid contamination, all surface brightness profiles shown in Fig. 4 are computed after applying a mask (see Fig. 10) to the images.
@@
With PSF effects removed, the galaxy’s outer light distribution now reliably represents the intrinsic stellar light of the galaxy and we can proceed to the analysis of the color and stellar mass density profiles.
The averaged color profile (shown in the lower panel of Fig. 4) of NGC 3486 shows a typical U-shaped profile (Bakos et al. 2008; Azzollini et al. 2008) with an inner red bulge ( 0.7) transitioning to a bluer disc with an average 0.3 at R = 100 and then reddening to 0.45 beyond R = 230.
We derive the 2D surface mass density map of NGC 3486 using the Mass-to-Light ratio (M/L)λ versus color relation prescribed in Roediger & Courteau (2015). Explicitly we use the Sloan data as reference as it is our deepest imaging dataset and the color. We assume a Chabrier Initial Mass Function (IMF; Chabrier 2003). The color image is created by subtracting the g and r surface brightness 2D maps and correcting of Galactic extinction using the following coefficients: = 0.08 and = 0.06 (Schlafly & Finkbeiner 2011). Inclination correction is included following the prescriptions in Trujillo et al. (2020). The resulting Cartesian and Polar maps of the stellar surface mass distribution of NGC 3486 are shown in the top left and top central panels of Fig. 5 respectively.
After constructing the stellar surface mass density map, we apply the mask generated in Sect. 4.1 to exclude sources not belonging to the galaxy. Then, to extract the radial profile, we use fixed elliptical apertures as in Sect. 4.2.1. The resulting stellar mass density profile is presented in orange in the bottom central panel of Fig. 5. The profile reveals an inner luminous component, followed by a gradual exponential decline out to R 200. Beyond this radius, the mass density profile steepens.
4.4 One dimensional edge detection:
We compute the numerical derivatives of the averaged stellar mass density profile of NGC 3486. The minimum of the second numerical derivative lies at 5888The error associated with the numerical derivative is half of the width of the bin we use to extract the profile.. To test the result, we fit a polynomial function to the average stellar mass density profile and we compute its analytical derivatives (its second derivative is shown as a blue deshed line in the bottom right panel of Fig. 5). The minimum obtained from the polynomial fit occurs at . Considering the uncertainty between these two approaches, we adopt a final = 205 5, indicated by the black dashed line. This radius corresponds to a mean surface stellar mass density of 1.3 0.1 , a = 25.60 0.01 mag/arcsec2 and a = 25.95 0.01 mag/arcsec2. We discuss the implications of these values in the following sections.

### v2

@@ 3 Target selection and data
We applied the methodology described above to the LIGHTS data of NGC 3486, an M33-like galaxy (i.e., with a similar stellar and total mass to M33) located at 13.6 Mpc (Kourkchi et al. 2020) and in a region of low Galactic extinction (Ag = 0.08, Ar = 0.06;
@@
NGC 3486 constitutes a case study; we aim to perform the analysis to the entire LIGHTS dataset.
@@
The LBC setup includes four CCDs, with each CCD covering 7.8 17.6 section of the sky and a total camera field of view (FoV) of 23 25.
@@
The data collection took place on April 3, 2021, under dark sky conditions. By using dithered exposures of 180 seconds each, with four arcminutes offsets, we obtained a total on-source integration time of 1.5 hours (see Zaritsky et al. 2024 for more details on observations).
The data reduction process followed a similar strategy to the one outlined in Trujillo et al. (2021) and developed in Zaritsky et al. (2024).
Our pipeline’s main steps involved deriving flat field images, background removal, astrometric and photometric calibrations, and mosaic co-addition.
@@
With this observational strategy and optimized data reduction we achieved a surface brightness limiting depth of 31.4 mag/arcsec2 and 30.5 mag/arcsec2 in the and Sloan bands, respectively, corresponding to a 3 fluctuation of the background in areas equivalent to 10 10 (Trujillo & Fliri 2016; Román et al. 2020).
@@ 4 Methods
The edge detection process for NGC 3486 follows a structured approach. We began by masking foreground and background sources to minimize contamination. Next, although LIGHTS data are background-subtracted, in some cases the subtraction was not precise enough to reliably represent the local surrounding background of the galaxy, leading to slight over- or underestimation. Therefore, we corrected for the local background. Then, we applied the point spread function (PSF) correction to account for the scattering of the galaxy’s light due to the effect of the PSF. After this, we proceeded to get the surface stellar mass density map of NGC 3486. Finally, we derived the location of the edge of NGC 3486 using the methodology described in Sect. 2. Figure 2 provides a schematic overview that serves as a reference for the methodology presented in the following sections.
@@ 4.1 Mask and local background treatment
To achieve accurate background subtraction, we first masked foreground and background sources. The mask was generated by combining the and band masks. The first step in creating the masks involves running NoiseChisel and Segment (Akhlaghi & Ichikawa 2015; Akhlaghi 2019). To improve the detection of contaminating sources, we applied unsharp masking, which removes the light from the galaxy and facilitates the identification of individual contamination sources. Unsharp masking involves smoothing the original image with a Gaussian filter corresponding to the FWHM of point-like sources and subtracting this smoothed image from the original. Finally, we manually inspected the deep color image of NGC 3486 (see Fig. 9) to mask any remaining undetected objects. The final mask applied to NGC 3486 is shown in Fig. 10 in Appendix C.
To compute the local background correction, once the data were masked, we manually placed circular apertures (each with a 20 radius) in specific regions approximately 6.5 from the galaxy center. These locations were selected to be relatively free of contamination, avoiding bright sources ( mag) and diffuse light brighter than mag/arcsec2. We adopted this conservative approach to ensure robust background estimation. We then computed the resistant (3-clipped) median of the unmasked background pixels within these regions. This process was performed separately for each filter.
@@ 4.2 Removing the effect of the PSF in the outermost regions of the galaxy
Light scattering from the PSF can significantly alter the observed properties of a galaxy’s outer disk by redistributing mostly light from the central regions, leading to an artificial brightening of the outskirts (Trujillo et al. 2001; Sandin 2014, 2015). The effect depends on the instrument and wavelength used (Duc et al. 2015; Croston et al. 2006; Mihos et al. 2015; Abraham & van Dokkum 2014). In this work, we mitigated PSF effects in the galaxy’s outermost regions by combining model-based convolution techniques (Trujillo & Fliri 2016; Rich et al. 2019; Peters et al. 2017b; Mart´ınez-Lombilla & Knapen 2019; Infante-Sainz et al. 2020) with direct image deconvolution methods, such as PyOperators (Karabal et al. 2017). The following sections detail the steps taken in this process.
@@ 4.2.1 Breaking the PSF-Stellar halo degeneracy
The problem: It can be challenging to distinguish a faint stellar halo present around galaxies from the light-broadening effects caused by the PSF. This confusion, which we term PSF-Stellar halo degeneracy, makes it difficult to tell if a galaxy’s outer light distribution is due to an extended stellar halo or merely an artifact of the PSF. To illustrate the problem, we generated two BrokenExponential models using the makeimage task in IMFIT, with the same PA, AR, and outer edge radius, . Model1 has an outer slope () that is 20 shallower than the one of Model2. In addition, Model2 is composed of a faint Exponential stellar halo component. We applied the extended band PSF of the LBC (Sedighi et al. 2025) to both models. The averaged stellar mass surface densities of Model1 and Model2 are illustrated in the left and central panels of Fig. 3, respectively, along with their components and the profile of the PSF. The parameters used to create the models are reported in Appendix D.
@@
The outer profile (R ¿ 200 pix) in both models can be interpreted as either only a PSF effect on the declining disk light or a faint stellar halo combined with the PSF. This uncertainty highlights the PSF-Stellar halo degeneracy, making it difficult to distinguish between PSF broadening and a different outer profile, , plus a faint stellar halo. This test can be extended to higher inclinations, which amplify PSF effects, emphasizing the need for accurate PSF deconvolution when analyzing faint outer structures in galaxies.
@@
In summary, the PSF redistributes photons, artificially broadening the outer regions of galaxies and smoothing the slope of the outer disk (). In general this depends on the shape of the PSF. This redistribution does not alter the location of the edge in the stellar mass density distribution. However, it is key to correct for the PSF to ensure accurate recovery of the intrinsic shape of the profile, particularly the outer slope ().
@@
Wiener-Hunt deconvolution to break the degeneracy: We used the extended PSF of the LIGHTS survey (Sedighi et al. 2025) to deconvolve the background-subtracted data by applying the Wiener-Hunt (Orieux et al. 2010) deconvolution algorithm. To mathematically frame the effect of the PSF, the observed data from the telescope can be expressed as a convolution of the “true” image, , with the PSF, , plus an additive noise term, :
@@
In an ideal, noise-free scenario, we can use Fourier transforms to work with these data components in frequency space. Specifically, the Fourier transforms, of and of , would allow us to isolate , the Fourier transform of the true image. From this, we could reconstruct by transforming back into spatial space. However, because acts as a low-pass filter, it tends to amplify high-frequency noise. Thus, a noise suppressor is needed. The deconvolution process using Wiener-Hunt filter was performed in the Fourier domain:
@@
where is the complex conjugate of and is the regularization parameter that controls noise suppression. In this work we used Wiener-Hunt deconvolution implemented in the Python package skimage666 skimage.restoration.unsupervised.Wiener, in which the regularization parameter is automatically estimated using a stochastic iterative process (Orieux et al. 2010).
@@
To visualize the effect of performing the Wiener-Hunt deconvolution on the data, we extracted the surface brightness profiles for NGC 3486 in both and bands with astscript-radial-profile and a bin size of 10777Larger bin sizes smooth the profiles by averaging the flux over wider elliptical annuli, thereby reducing point-to-point fluctuations. In contrast, smaller bin sizes increase the sensitivity to small-scale variations, leading to more oscillatory behavior in the derivative profiles. We verified that the choice of step size does not affect the determination of the edge location.. We extracted the light profiles of the galaxy using AR = and a PA of (clockwise from the X axis). This choice of parameters is justified in the Appendix E and compared with previous estimates. To maintain consistency, we applied the same mask used for sky background estimation (see Fig. 10). The errors on these profiles were calculated by combining the uncertainties of each elliptical annulus with the one of the local sky estimate, as is outlined in Golini et al. (2024).
@@
In Fig. 4, we show in green the resulting and surface brightness profiles prior to deconvolution, which exhibit an inner component extending to about 40, followed by a exponential decline. At around 200, a slope change occurs, with the profile steepening in both bands, marking the edge of the visible disk seen in Fig. 9.
@@
The Wiener-Hunt deconvolved surface brightness profiles in the and bands are also shown in Fig. 4. Compared to the original (background-subtracted) profiles, the deconvolved profiles exhibit a steeper outer slope beyond R = 190. We visually define this transition point as . This effect is due to the redistribution of light from the galaxy’s outskirts to the inner regions, compensating for the PSF-induced broadening in the original data. For instance, post-deconvolution, the inner region (R ¡ 20) appears 0.1 magnitude brighter, while outer regions (R ¿ ″) are up to 1 magnitude fainter. The edge location is preserved.
@@
The Wiener-Hunt filter is a widely used deconvolution technique, but like other deconvolution methods such as Richardson-Lucy (Richardson 1972; Lucy 1974) or ones based on spatial regularizations such as the assumption of sparsity in a wavelet domain (Starck & Murtagh 1994; Starck et al. 2006), it introduces image smoothness, correlates pixel values, and alters background statistics, especially in regions of LSB (Murli et al. 1999; Orieux et al. 2010; Ramos et al. 2020; Yoo & Ahn 2014).
@@
4.2.2 Creation of a model of the galaxy not affected by PSF
To remove the PSF effect from the NGC 3486 data, while preserving the noise structure of the LIGHTS image, we proceeded as follows. First, we modeled the averaged surface brightness profile of NGC 3486 in both bands after Wiener-Hunt deconvolution (blue profiles in Fig. 4) using a custom implementation of the IMFIT Exponential function in Python.
We modeled the Wiener-Hunt deconvolved profile in the range between (vertical dashed purple lines in Fig. 4), which is the point at which the PSF begins to significantly affect the outermost regions of the light profiles in the and filters (i.e., 25 mag/arcsec2), to the radius where the uncertainty reaches 0.1 mag. This results in an outer exponential scale length () of 18.8 in both filters. Second, we generated a 2D image of this Exponential model using makeimage. When constructing the 2D outer disk component, we adopted the PA and inclination parameters obtained in Appendix E. We did not include a bulge or stellar halo component to avoid imposing a priori assumptions. Stellar halos often exhibit asymmetries, spatial offsets, or orientations misaligned with the disk, making them difficult to model reliably. Our final model of the galaxy, not affected by the PSF, consists of two parts: an inner region (R ¡ ) represented by the Wiener-Hunt deconvolved image, which enhances spiral structures that would otherwise be difficult to model, and an outer region (R ¿ ) that transitions to an exponential declining profile (shown in pink in Fig. 4).
@@ 4.2.3 Subtraction of the model and residuals addition: PSF deconvolved data
After generating the two deconvolved models (one for and one for the band) of NGC 3486, we employed the astconvolve tool in Gnuastro to convolve these models with the LIGHTS’ PSF. To isolate structural components beyond the PSF effect, we subtracted these convolved images from the original galaxy data using the astarithmetic tool in Gnuastro, resulting in a “residuals” image. These residuals highlight the galaxy’s asymmetric faint features in the outer region of the galaxy that are not accounted for in the symmetric model.
Following this, we added the residuals image back into the model created in Sect. 4.2.2, resulting in a refined image with PSF effects effectively removed. This stage of analysis ensures that any faint residual light detected in the galaxy’s outer regions corresponds to real, low-luminosity stellar structures rather than artifacts introduced by the PSF.
The surface brightness profiles, after PSF correction, extracted using fixed ellipses, are shown in orange in Fig. 4. In both bands, the galaxy light distribution after PSF removal shows a steeper decline, up to one magnitude fainter, starting at R ″. We remark that, to avoid contamination, all surface brightness profiles shown in Fig. 4 were computed after applying a mask (see Fig. 10) to the images.
@@
With PSF effects removed, the galaxy’s outer light distribution now reliably represents the intrinsic stellar light of the galaxy and we can proceed to the analysis of the color and stellar mass density profiles. The averaged color profile (shown in the lower panel of Fig. 4) of NGC 3486 shows a typical U-shaped profile (Bakos et al. 2008; Azzollini et al. 2008) with an inner red bulge ( 0.7) transitioning to a bluer disk with an average 0.3 at R = 100 and then reddening to 0.45 beyond R = 230.
We derived the 2D surface mass density map of NGC 3486 using the mass-to-light ratio (M/L)λ versus color relation prescribed in Roediger & Courteau (2015). Explicitly we used the Sloan data as reference as it is our deepest imaging dataset and the color. We assumed a Chabrier initial mass function (IMF; Chabrier 2003). The color image was created by subtracting the g and r surface brightness 2D maps and correcting for Galactic extinction using the following coefficients: = 0.08 and = 0.06 (Schlafly & Finkbeiner 2011). Inclination correction was included following the prescriptions in Trujillo et al. (2020). The resulting Cartesian and polar maps of the stellar surface mass distribution of NGC 3486 are shown in the top left and top central panels of Fig. 5, respectively.
After constructing the stellar surface mass density map, we applied the mask generated in Sect. 4.1 to exclude sources not belonging to the galaxy. Then, to extract the radial profile, we used fixed elliptical apertures as in Sect. 4.2.1. The resulting stellar mass density profile is presented in orange in the bottom central panel of Fig. 5. The profile reveals an inner luminous component, followed by a gradual exponential decline out to R 200. Beyond this radius, the mass density profile steepens.
4.4 One-dimensional edge detection:
We computed the numerical derivatives of the averaged stellar mass density profile of NGC 3486. The minimum of the second numerical derivative lies at 5888The error associated with the numerical derivative is half of the width of the bin we use to extract the profile.. To test the result, we fit a polynomial function to the average stellar mass density profile and we computed its analytical derivatives (its second derivative is shown as a dashed blue line in the bottom right panel of Fig. 5). The minimum obtained from the polynomial fit occurs at . Considering the uncertainty between these two approaches, we adopted a final = 205 5, indicated by the dashed black line. This radius corresponds to a mean surface stellar mass density of 1.3 0.1 , a = 25.60 0.01 mag/arcsec2 and a = 25.95 0.01 mag/arcsec2. We discuss the implications of these values in the following sections.